# Sparsity Outperforms Low-Rank Projections in Few-Shot Adaptation

### v1

@@ 1 Introduction
The performance of deep learning (DL) models is strongly influenced by their size, with larger architectures yielding superior results. For instance, the ViT-H/14 backbone outperforms ViT-B/32 by a significant margin across multiple benchmarks [6]. Despite recent architectural advancements, the best models remain highly data-dependent and require large-scale labeled datasets to achieve reliable generalization. While techniques such as data augmentation and regularization help mitigate overfitting in low-data scenarios, they are insufficient to fully resolve it. In contrast, the human brain has more than billion neurons. Each neuron forms, on average, only about synaptic connections, which results in an estimated total of to trillion synapses. Despite its tremendous capacity, the brain exhibits a remarkable ability to grasp new concepts from minimal exposure. This one-shot or few-shot learning capability might be partly attributed to the effective use of sparsity [17, 40]. By selectively activating a fraction of connections at any given time, the brain optimizes both energy consumption and computational efficiency.
@@
Among the prominent DL models, vision-language models (VLMs) learn visual concepts from natural language descriptions. In particular, the Contrastive Language-Image Pretraining (CLIP) [33] model is pretrained on large-scale datasets containing hundreds of millions of image-text pairs to align visual and textual embeddings within a shared semantic space. After training, VLMs show strong zero-shot capabilities across diverse visual recognition tasks [33]. Moreover, fine-tuning VLMs yields state-of-the-art performance in various domain adaptation scenarios [26, 39]. Despite their zero-shot capabilities, adapting VLMs to new tasks with very limited labeled datasets remains challenging due to pronounced overfitting and catastrophic forgetting problems. Furthermore, training such models requires vast resources, including significant GPU memory for storing millions of parameters, gradients, activations, and optimizer states. These challenges underscore the need for cost-effective optimization strategies that address overfitting.
@@
To address the limitations of full VLM fine-tuning in few-shot learning, state-of-the-art methods leverage parameter-efficient training (PET) [16, 15]. This strategy consists of optimizing a small number of parameters compared to the full set of weights. Additive PET methods freeze the initial weights and attach trainable and unmergeable parameters to the model. Within this category, we find the adapter and the prompt tuning variants. The prevailing methods in few-shot VLMs have mainly focused on this category. Adjustable weights can be incorporated at various levels, either by embedding them directly into the input space as visual [18] or textual prompts [49], or by integrating adapters throughout the network architecture [51].
@@
To mitigate overfitting in few-shot adaptation and preserve the prior knowledge of VLMs, adapter-based methods typically employ three techniques. First, these small trainable modules are only integrated into the last layer of VLMs. Second, they use regularization techniques to penalize the divergence between the initial zero-shot class prototypes and the learnable class prototypes. Finally, they initialize the learnable class prototypes into a reliable region using the zero-shot prototypes. As a result, these methods have shown better results than the prompt tuning approaches [45], which often introduce computational overhead [46]. Despite these efforts, adapter-based methods are still prone to considerable overfitting [34]. Their success depends heavily on carefully tuning hyperparameters whose optimal values vary significantly across different tasks [34]. This sensitivity leads to performance degradation that undermines the generalizability of these approaches. It is worth noting that tuning these hyperparameters requires extensive grid searches using a validation set, which is not available in real-world few-shot scenarios.
@@
Recently, low-rank reparameterization methods have gained popularity for domain adaptation. These approaches reparameterize the initial weights with supplementary trainable and mergeable low-rank matrices. Low-Rank Adaptation (LoRA) [16] was the first to apply this strategy. It reformulates weight matrices as , where and are low-rank matrices. As a key advantage, this technique helps reduce overfitting and catastrophic forgetting because of two properties. First, LoRA enables fine-grained parallel adaptation by learning small perturbations to the frozen weights [13]. Second, the effective initialization of and helps in preserving prior knowledge [29, 44].
@@
Building on these advantages, recent studies [46, 24] show that low-rank reparameterization methods achieve state-of-the-art performance in few-shot adaptation of VLMs. However, these methods also have notable limitations. Compelling evidence suggests that restricting weight updates to low-rank subspaces reduces model expressiveness [48, 12, 23, 31, 37, 11, 4]. In this work, we argue that the low-rank projection technique exhibits severe overfitting in few-shot adaptation. Our systematic study of LoRA confirms this issue (see Fig. 1). Test-set accuracy initially improves but then declines sharply. At convergence, when the loss approaches zero, performance deteriorates significantly compared to its peak on test-set data. Moreover, we can see that the optimal number of training iterations varies widely across datasets. However, tuning hyperparameters on a validation set contradicts the principles of few-shot learning.
@@
The challenges of low-rank reparameterization methods extend beyond fixing the number of training iterations. Their performance is also highly sensitive to another hyperparameter: the rank of the low-rank projection. This sensitivity is illustrated in Fig. 2, which illustrates LoRA’s performance for different ranks on three datasets (DTD, Oxford Pets, UCF101) in a 1-shot setting. Each dataset reveals a distinct best rank and higher ranks sometimes lead to an early spike in accuracy but then degrade as training continues. Further experiments examining these behaviors are presented in Appendix F of the supplementary material.
@@
Since rank is a discrete value that cannot be lower than 1, the low-rank methods lack the flexibility to adjust very few parameters. While setting the rank to a very low value (e.g., 1 or 2) can help mitigate overfitting, this choice might reduce the model’s learning capacity. This limitation arises because the trainable parameters in low-rank methods remain fixed throughout training, potentially leading to underfitting the VLM capacity. Indeed, rank selection involves a strong trade-off between underfitting and overfitting (see Fig. 2). For instance, in the 1-shot DTD setting, LoRA at rank 5 briefly exceeds rank 2 but then drops below it as training continues. Overall, the test accuracy of LoRA exhibits oscillations and varies unpredictably w.r.t. rank and the number of training iterations. This instability makes it difficult to reliably tune these interdependent hyperparameters without access to a validation set.
@@
In this paper, we deviate from the current long-standing trend of low-rank adaptation and introduce two new paradigms that make sparsity work in few-shot settings. Our approach is a simple yet effective optimization strategy, referred to as Sparse Optimization (SO). As a first paradigm, we advocate for local sparsity and global density. Unlike the conventional low-rank approaches that update a fixed set of parameters, SO dynamically updates a minimal subset of original connections at each step. Specifically, high sparsity is enforced in both the gradient and moment updates while allowing the sparsity support to evolve dynamically throughout training. This paradigm ensures that, at any given iteration, only a minimal subset of parameters is updated to reduce the local learning capacity and prevent rapid overfitting. By dynamically changing the sparsity support, the model retains its overall expressiveness and ensures that different parameters contribute to learning over time.
@@
As a second paradigm, we advocate for local randomness and global importance. Similar to Adam [19], we leverage the gradient and the first and second moments. However, we sparsify these components in a way that addresses overfitting and ensures efficiency in terms of memory consumption. More precisely, we sparsify the gradient randomly and the first moment by importance. The gradient captures local and iteration-specific information. The first moment aggregates the gradients over the whole path and, thus, reflects long-term parameter importance. Random pruning of the gradient prevents the model from relying too much on short-term and local high-magnitude updates. The importance-based selection of the first moment ensures that connections with long-term significance are updated. By emphasizing local randomness and global importance, our strategy avoids severe overfitting.
@@
For a given density rate , we retain elements for both the gradient and moment updates at each iteration. First, we randomly select gradient values from the full set of parameter updates. Next, we compute the temporary first and second moments, which can contain up to values in the worst-case scenario where no overlap exists between the selected gradients and the previous first moment values. We rank the first moment values by importance using their magnitude. Only the top values are retained for the first moment. We then use the same indices to select the corresponding values for the second moment. This ensures alignment between the sparsified first and second moments. The selection strategy preserves the most relevant long-term updates while introducing randomness in the local updates.
@@
Contributions. (1) A novel SO framework is proposed that sparsifies both gradients and moments to improve generalization and mitigate overfitting in few-shot learning. Our approach differs from conventional low-rank reparameterization by focusing on sparsity. (2) We introduce the local sparsity and global density paradigm, where a dynamically evolving sparsity support updates only a minimal subset of model parameters at each iteration. This strategy preserves the overall model expressiveness while reducing its risk of overfitting. (3) We propose the local randomness and global importance paradigm, which selectively sparsifies the gradient using random selection while pruning the first moment based on importance. This selection strategy preserves the most relevant long-term updates while introducing local randomness to mitigate overfitting. (4) Our comprehensive empirical study on 11 diverse datasets compares our method to state-of-the-art low-rank approaches. Results indicate that SO can significantly improve few-shot generalization while achieving memory efficiency.
@@ 3 Proposed Sparse Optimization Method
0 | CLIP (ICML ’21) | 66.7 | 62.6 | 24.7 | 47.5 | 65.3 | 86.1 | 89.1 | 71.4 | 92.9 | 43.6 | 66.7 | 65.1 |
1 | Adam (ICLR ’15) | 0.1 | 0.6 | 1.0 | 24.3 | 0.5 | 1.0 | 2.9 | 1.4 | 1.0 | 5.6 | 1.7 | 3.6 |
@@
SO | ✗ | ✓ | ✗ | ✓ | 38.6 | 87.7 | 78.9 | 85.3 | 92.4 | 95.1 | 95.5 | 66.4 | 83.4 | 80.4 |
Algorithm. Our algorithm is provided in Appendix B due to space limitations. We finetune CLIP until convergence using our optimization technique. To promote efficiency, we retain the gradient sparsity support fixed for iterations before reselecting the trainable parameters. Overall, our approach introduces two hyperparameters: the density ratio , which controls the fraction of retained gradient entries, and the update interval , which determines how frequently the sparsity support is refreshed. A smaller value improves generalization but results in slower convergence. In contrast, a larger accelerates optimization at the cost of reduced performance. Similarly, reducing leads to better results by enabling more frequent updates to the sparsity suport, although increasing speeds up convergence.
@@ G.3 Results
4 | Importance-Based Gradient Pruning | |||||||
@@
8 | Importance-Based Gradient Pruning | |||||||
@@
16 | Importance-Based Gradient Pruning | |||||||
@@
4 | Importance-Based Gradient Pruning | ||||||
@@
8 | Importance-Based Gradient Pruning | ||||||
@@
16 | Importance-Based Gradient Pruning | ||||||
SO () | |||||||
SO () | |||||||
SO () | |||||||
SO () | |||||||
SO () | |||||||
Random Gradient Pruning | |||||||
SO () | |||||||
SO () | |||||||
SO () | |||||||
SO () | |||||||
SO () |
4 | Importance-Based Gradient Pruning | ||||||
@@
8 | Importance-Based Gradient Pruning | ||||||
@@
16 | Importance-Based Gradient Pruning | ||||||
@@
SO () |

### v2

@@ 1 Introduction
The performance of deep learning (DL) models is strongly influenced by their size, with larger architectures yielding superior results. For instance, the ViT-H/14 backbone outperforms ViT-B/32 by a significant margin across multiple benchmarks [6]. Despite recent architectural advancements, the best models remain highly data-dependent and require large-scale labeled datasets to achieve reliable generalization. While techniques such as data augmentation and regularization help mitigate overfitting in low-data scenarios, they are insufficient to fully resolve it [52]. In contrast, the human brain has more than billion neurons. Each neuron forms, on average, only about synaptic connections, which results in an estimated total of to trillion synapses. Despite its tremendous capacity, the brain exhibits a remarkable ability to grasp new concepts from minimal exposure. This one-shot or few-shot learning capability might be partly attributed to the effective use of sparsity [17]. By selectively activating a fraction of connections at any given time, the brain optimizes both energy consumption and computational efficiency.
@@
Among the prominent DL models, vision-language models (VLMs) learn visual concepts from natural language descriptions. In particular, the contrastive language-image pretraining (CLIP) [33] model is pretrained on large-scale datasets containing hundreds of millions of image-text pairs to align visual and textual embeddings within a shared semantic space. After training, VLMs show strong zero-shot capabilities across diverse visual recognition tasks [33]. Moreover, fine-tuning VLMs yields state-of-the-art performance in various domain adaptation scenarios [26, 39]. Despite their zero-shot capabilities, adapting VLMs to new tasks with small labeled datasets remains challenging due to pronounced overfitting and catastrophic forgetting problems. Furthermore, training such models requires vast resources, including significant GPU memory for storing millions of parameters, gradients, activations, and optimizer states. These challenges underscore the need for cost-effective optimization strategies that address overfitting.
@@
To address the limitations of full VLM fine-tuning in few-shot learning, state-of-the-art methods leverage parameter-efficient training (PET) [16, 15, 20]. This strategy consists of optimizing a small number of parameters compared to the full set of weights. Additive PET methods freeze the initial weights and attach trainable and unmergeable parameters to the model. Within this category, we find the adapter [53] and the prompt tuning variants [50]. The prevailing methods in few-shot VLMs have mainly focused on this category. Adjustable weights can be incorporated at various levels, either by embedding them directly into the input space as visual [18] or textual prompts [49], or by integrating adapters throughout the network architecture [51].
@@
To mitigate overfitting in few-shot adaptation and preserve the prior knowledge of VLMs, adapter-based methods typically employ three techniques. First, these small trainable modules are only integrated into the last layer of VLMs. Second, they use regularization techniques to penalize the divergence between the initial zero-shot class prototypes and the learnable class prototypes. Finally, they initialize the learnable class prototypes into a reliable region using the zero-shot prototypes. As a result, these methods have shown better results than the prompt tuning approaches [45], which often introduce computational overhead [46]. Despite these efforts, adapter-based methods are still prone to considerable overfitting [34]. Their success depends heavily on carefully tuning hyperparameters whose optimal values vary significantly across different tasks [34]. This sensitivity leads to performance degradation that undermines the generalizability of these approaches. It is worth noting that tuning these hyperparameters requires extensive grid searching with a validation set, which is not available in real-world few-shot scenarios.
@@
Recently, low-rank reparametrization [16, 25, 54] has gained popularity for domain adaptation. These approaches reparameterize the initial weights with supplementary trainable and mergeable low-rank matrices. Low-rank adaptation (LoRA) [16] was the first to apply this strategy. It reformulates weight matrices as , where and are low-rank matrices. As a key advantage, this technique helps reduce overfitting and catastrophic forgetting because of two properties. First, LoRA enables fine-grained parallel adaptation by learning small perturbations to the frozen weights [13]. Second, the effective initialization of and helps in preserving prior knowledge [29, 44].
@@
Building on these advantages, recent studies [46, 24] show that low-rank reparameterization methods achieve state-of-the-art performance in few-shot adaptation of VLMs. However, these methods also have notable limitations. Compelling evidence suggests that restricting weight updates to low-rank subspaces reduces model expressiveness [48, 23, 37]. In this work, we argue that the low-rank projection technique exhibits severe overfitting in few-shot adaptation. Our systematic study of LoRA confirms this issue (see Fig. 1). Test-set accuracy initially improves but then declines sharply. At convergence, when the loss approaches zero, performance deteriorates significantly compared to its peak on test-set data. Moreover, we can see that the optimal number of training iterations varies widely across datasets. However, tuning hyperparameters, including the number of iterations for early stopping, on a validation set contradicts the principles of few-shot learning.
@@
The challenges of low-rank reparameterization methods extend beyond fixing the number of training iterations. Their performance is also highly sensitive to another hyperparameter: the rank of the low-rank projection. This sensitivity is shown in Fig. 2, which illustrates LoRA’s performance for different ranks on three datasets (DTD, Oxford Pets, UCF101) in a 1-shot setting. Each dataset reveals a distinct best rank, and higher ranks sometimes lead to an early spike in accuracy but then degrade as training continues. Further experiments examining these behaviors are presented in Appendix F of the supplementary material.
@@
Since rank is a discrete value that cannot be lower than 1, the low-rank methods lack the flexibility to adjust very few parameters. While setting the rank to a very low value (e.g., 1 or 2) can help mitigate overfitting, this choice might reduce the model’s learning capacity. This limitation arises because the trainable parameters in low-rank methods remain fixed throughout training, potentially leading to underfitting the VLM capacity. Indeed, rank selection involves a trade-off between underfitting and overfitting (see Fig. 2). For instance, in the 1-shot DTD setting, LoRA at rank 5 briefly exceeds rank 2 but then drops below it as training continues. Overall, the test accuracy of LoRA exhibits oscillations and varies unpredictably w.r.t. rank and the number of training iterations. This instability makes it difficult to reliably tune these interdependent hyperparameters without access to a validation set.
@@
In this paper, we deviate from the current long-standing trend of low-rank adaptation and introduce two new paradigms that make sparsity work in few-shot settings. Our approach is a simple yet effective optimization strategy, referred to as sparse optimization (SO), and aligns with these paradigms. As a first paradigm, we advocate for local sparsity and global density. Unlike the conventional low-rank approaches that update a fixed set of parameters, SO dynamically updates a minimal subset of original connections at each step. Specifically, high sparsity is enforced in both the gradient and moment updates while allowing the sparsity support to evolve dynamically throughout training. This paradigm ensures that, at any given iteration, only a minimal subset of parameters is updated to reduce the local learning capacity and prevent rapid overfitting. By dynamically changing the sparsity support, the model retains its overall expressiveness and ensures that different parameters contribute to learning over time.
@@
As a second paradigm, we advocate for local randomness and global importance. Similar to Adam [19], we leverage the gradient and the first and second moments. However, we sparsify these components in a way that addresses overfitting and ensures efficiency in terms of memory consumption. In particular, we sparsify the gradient randomly and the first moment by importance. The gradient captures local and iteration-specific information. The first moment aggregates the gradients over the whole path and, thus, reflects long-term parameter importance. Random pruning of the gradient prevents the model from relying too much on short-term and local high-magnitude updates. The importance-based selection of the first moment ensures that connections with long-term significance are updated. By emphasizing local randomness and global importance, our strategy avoids severe overfitting.
@@
Concretely instantiating this paradigm, we fix a density rate and retain elements for both the gradient and moment updates at each iteration. First, we randomly select gradient values from the full set of parameter updates. Next, we compute the temporary first and second moments, which can contain up to values in the worst-case scenario where no overlap exists between the selected gradients and the previous first moment values. We rank the first moment values by importance using their magnitude. Only the top values are retained for the first moment. We then use the same indices to select the corresponding values for the second moment. This ensures alignment between the sparsified first and second moments. The selection strategy preserves the most relevant long-term updates while introducing randomness in the local updates.
@@
Our contributions. (1) A novel framework is proposed that sparsifies both gradients and moments to improve generalization and mitigate overfitting in few-shot learning. Our approach differs from conventional low-rank reparameterization by focusing on sparsity. (2) We introduce the local sparsity and global density paradigm, where a dynamically evolving sparsity support updates only a minimal subset of model parameters at each iteration. This strategy preserves the overall model expressiveness while reducing its risk of overfitting. (3) We propose the local randomness and global importance paradigm, which selectively sparsifies the gradient using random selection while pruning the first moment based on importance. This selection strategy preserves the most relevant long-term updates while introducing local randomness to mitigate overfitting. (4) Our comprehensive empirical study on 11 diverse datasets compares our method to state-of-the-art low-rank approaches. Results indicate that SO can significantly improve few-shot generalization while achieving memory efficiency.
@@ 3 Proposed Sparse Optimization Method
| 0 | CLIP (ICML ’21) | 66.7 | 62.6 | 24.7 | 47.5 | 65.3 | 86.1 | 89.1 | 71.4 | 92.9 | 43.6 | 66.7 | 65.1 |
| 1 | Adam (ICLR ’15) | 0.1 | 0.6 | 1.0 | 24.3 | 0.5 | 1.0 | 2.9 | 1.4 | 1.0 | 5.6 | 1.7 | 3.6 |
@@
| SO | ✗ | ✓ | ✗ | ✓ | 38.6 | 87.7 | 78.9 | 85.3 | 92.4 | 95.1 | 95.5 | 66.4 | 83.4 | 80.4 | |
Algorithm. Our algorithm is provided in Appendix B due to space limitations. We finetune CLIP until convergence using our optimization technique. To promote efficiency, we retain the gradient sparsity support fixed for iterations before reselecting the trainable parameters. Overall, our approach introduces two hyperparameters: the density ratio , which controls the fraction of retained gradient entries, and the update interval , which determines how frequently the sparsity support is refreshed. A smaller value improves generalization but results in slower convergence. In contrast, a larger accelerates optimization at the cost of reduced performance. Similarly, reducing leads to better results by enabling more frequent updates to the sparsity support, although increasing speeds up convergence.
@@ 5 Conclusion
Acknowledgement
This work was supported by the Natural Sciences and Engineering Research Council of Canada (NSERC), and the Digital Research Alliance of Canada.
@@ G.3 Results
| 4 | Importance-Based Gradient Pruning | |||||||
@@
| 8 | Importance-Based Gradient Pruning | |||||||
@@
| 16 | Importance-Based Gradient Pruning | |||||||
@@
| 4 | Importance-Based Gradient Pruning | ||||||
@@
| 8 | Importance-Based Gradient Pruning | ||||||
@@
| 16 | Importance-Based Gradient Pruning | ||||||
@@
| 4 | Importance-Based Gradient Pruning | ||||||
@@
| 8 | Importance-Based Gradient Pruning | ||||||
@@
| SO () | |||||||
| 16 | Importance-Based Gradient Pruning | ||||||
| SO () | |||||||
| SO () | |||||||
| SO () | |||||||
| SO () | |||||||
| SO () | |||||||
| Random Gradient Pruning | |||||||
| SO () | |||||||
| SO () | |||||||
| SO () | |||||||
| SO () | |||||||
| SO () | |||||||